Main INDEX
Monthly INDEX
PREV
NEXT
User name jroche
Log entry time 23:56:07 on July18,2004
Entry number 128910
keyword=swing shift summary
Highlights
- Multiple target ioc reboot
- change HWP state at 19:38 and then at 23:54
- a profile scan run
- a lot of good data
Time line
- 16:27 : we are ready for beam (speta up). MCC is still having problem with on of the RF zone
- 17:37 :MCC is sending us 38 uA
- 18:37 Achemi changes the HV for the lumi detectors between 2952 and 2953, done at 18:58
- 19:00 Kirsten reboot the target IOC
- 19:38 HWP removed
- 20:00 yet another target reboot
- 22:27 started a profile scan
Run taken during this shift
- 2950,2051 : pedestal
- 2952,2953 : production hwp IN
- 2954,2955,2956 : production hwp OUT
- 2957 : profile scan
- 2958 : production OUT
- 2959 : production IN